Calculate the period of a simple pendulum of length 0.5m in a place were the acceleration due to gravity is 10ms-2
1
Expert's answer
2020-04-06T08:41:55-0400

The period of simple pendulum

"T=2\pi \sqrt{\frac{l}{g}}=2\pi \sqrt{\frac{0.5}{10}}=1.4\:\rm s."